OpenStudy (anonymous):

~~~~~~~~~~~~~`FAN AND MEDAL`~~~~~~~~~~ What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

first find the slope second use y-y1 = m(x-x1) and plug the points. m = slope then put the equation in to slope intercept form y= mx+b then set y to zero and solve for x

OpenStudy (the_fizicx99):

To get an intercept just multiply the y by 0 so you just have your x

OpenStudy (anonymous):

I know this sounds so stupid but I dont know which one is the x an which is the y

OpenStudy (anonymous):